# How Far Are We to GPT-4V? Closing the Gap to Commercial Multimodal Models with Open-Source Suites

### v1

@@ Abstract
In this report, we introduce InternVL 1.5, an open-source multimodal large language model (MLLM) to bridge the capability gap between open-source and proprietary commercial models in multimodal understanding. We introduce three simple improvements: (1) Strong Vision Encoder: we explored a continuous learning strategy for the large-scale vision foundation model—InternViT-6B, boosting its visual understanding capabilities, and making it can be transferred and reused in different LLMs. (2) Dynamic High-Resolution: we divide images into tiles ranging from 1 to 40 of 448$\times$448 pixels according to the aspect ratio and resolution of the input images, which supports up to 4K resolution input. (3) High-Quality Bilingual Dataset: we carefully collected a high-quality bilingual dataset that covers common scenes, document images, and annotated them with English and Chinese question-answer pairs, significantly enhancing performance in OCR- and Chinese-related tasks. We evaluate InternVL 1.5 through a series of benchmarks and comparative studies. Compared to both open-source and proprietary models, InternVL 1.5 shows competitive performance, achieving state-of-the-art results in 8 of 18 benchmarks.
@@ 1 Introduction
We evaluated InternVL 1.5 on 18 representative multimodal benchmarks, which are categorized into four specific groups: OCR-related, general multimodal, mathematical, and multi-turn conversation benchmarks. Compared to both open-source and proprietary models, InternVL 1.5 shows competitive performance, achieving state-of-the-art results in 8 of 18 benchmarks. Notably, as shown in Figure 1, it even surpasses leading proprietary models like Grok-1.5V [120], GPT-4V [83], Claude-3 Opus [5], and Gemini Pro 1.5 [88] in four specific benchmarks, particularly in OCR-related datasets such as TextVQA [96], ChartQA [77], and DocVQA [78]. This evaluation indicates that InternVL 1.5 has effectively narrowed the gap between open-source models and leading commercial models. We hope that our approach and open-source model weights can contribute to the development of the multimodal community.
@@ 2.2 Open-Source MLLMs
The development of open-source MLLMs [133, 134, 119, 98, 4, 43, 47, 52, 49, 20, 65, 66, 105, 113, 115, 15] has significantly influenced the AGI landscape by integrating and enhancing capabilities in processing both visual and textual data. Over the past year, many open-source MLLMs have become well-known, including the LLaVA series [59, 58, 60], MiniGPT-4 [137], VisionLLM [111], Qwen-VL [7], CogVLM [112], Shikra [17], and others [114, 86, 20, 25]. However, these models are typically trained on images with small, fixed resolutions such as 336$\times$336, 448$\times$448, or 490$\times$490, which leads to sub-optimal performance on images with unusual aspect ratios or document data. To address this issue, many approaches have been explored for training on high-resolution images. Currently, there are two common technical routes: one involves designing a dual-branch image encoder [32, 73, 116, 72], and the other involves dividing a high-resolution image into many low-resolution tiles [33, 46, 52, 53, 60, 64, 121, 122]. Despite these explorations in high-resolution training, these open-source models still exhibit significant gaps in understanding documents, charts, and infographics, as well as recognizing scene texts, compared to leading commercial models.
@@ 3.4 High-Quality Bilingual Dataset
Pre-training Dataset. The pre-training dataset utilized in our InternVL 1.5 encompasses a diverse range of publicly accessible sources. We provide an overview of these datasets in Table 1(a). These datasets span multiple tasks, including captioning, which predominantly uses datasets such as Laion-EN [89], Laion-ZH [89], COYO [12], and GRIT [86], constituting 53.9% of the total data. Detection and grounding tasks utilize datasets like Objects365 [93], GRIT [86], and All-Seeing [114], making up 5.2%. For OCR tasks, we utilized large-scale datasets such as Wukong-OCR, LaionCOCO-OCR, and Common Crawl PDFs, which constitute 32.0% of our data. These datasets were constructed using PaddleOCR [48] to perform OCR on Chinese images from Wukong [29] and on English images from Laion-COCO [90]. Smaller OCR datasets include MMC-Inst [57], LSVT [100], ST-VQA [11], RCTW-17 [94], and others, accounting for 8.9% of the data, which focus on more specific or constrained OCR challenges. This diverse dataset assembly ensures robust model pre-training of InternVL, catering to varied linguistic and visual elements across tasks.
@@
XComposer2 [25] 8B 55.7 51.0
Yi-VL-34B [125] 34B 54.2 48.6
Monkey-Chat [102] 10B 53.4 47.0
@@ 4.2.1 Quantitative Results on 18 Benchmarks
Figure 7: Examples for General QA. Blue highlights the standout answers.
@@ 4.3.1 Qualitative Results on Different Scenes
General QA. To compare the general capabilities of InternVL 1.5 and GPT-4V, we first conducted an experiment involving a simple user query with an image requiring general knowledge. As shown on the left side of Figure 7, both InternVL 1.5 and GPT-4V respond accurately to the query, showcasing their proficiency in general topics. As shown on the right side of Figure 7, GPT-4V may excessively refuse to answer some questions due to its involvement in personal privacy.
@@
   Refer to caption User: What is the seventh piece of advice in the image? GPT-4V: The image shows a signboard with Chinese characters on it. From the text, we can deduce that the sign is likely for a business. The sign features a design that includes a martini glass with a flamingo, indicating that the establishment might serve drinks. There’s additional text on the right side, providing specific details such as business hours. It appears to be a nighttime business, likely a bar or restaurant, operating from 16:30 to 22:30. The other text provides further information about the services or specialties of the place. Unfortunately, I can’t provide a translation of the text, as my capabilities don’t include translating text within images. InternVL 1.5: The image shows a signboard with Chinese text, indicating a business establishment. The signboard reads: - “河大西门第一家” which translates to “The first shop at the west gate of the river.” - “正宗 老字号 五香味” which means “Genuine, old brand, five-spice flavor.” - “曹记焖烤鸡腿” which translates to “Cao’s roasted and baked chicken legs.” - “营业时间 16:30-22:30” indicating the business hours from 4:30 PM to 10:30 PM. - “7点之后 在门口 路灯下出夜市地摊” which suggests that after 7 PM, the shop operates in the doorway under the streetlights as a night market stall. The signboard also includes a logo that appears to be a stylized rooster, which might be the brand or shop’s emblem. The overall style of the signboard is quite traditional and straightforward, commonly seen in Chinese street markets or small businesses.
@@
Figure 8: Examples for OCR-related QA. Standout answers are highlighted in blue, while incorrect responses are highlighted in red.
@@
Figure 11: Examples for Object Localization. Blue highlights the standout answers.
@@
Object Localization. Evaluating machine learning models for their proficiency in Object Localization tasks is essential, especially in applications requiring precise spatial awareness. In our comparative analysis, the performance of the InternVL 1.5 model was juxtaposed with GPT-4V, focusing on their ability to accurately detect and localize objects within various environments. Our assessments ranged from simple object recognition in cluttered scenes to complex scenarios involving dynamic interactions among multiple entities. As illustrated in Figure  LABEL:, the results demonstrate that InternVL 1.5 not only localized objects with high accuracy but also exhibited a comparable understanding of spatial relationships, matching the performance of GPT-4V.
@@ 5 Conclusion
This work introduced InternVL 1.5, an open-source multimodal large language model (MLLM) designed to narrow the performance gap between open-source and proprietary commercial models in multimodal understanding. By integrating a strong vision encoder with continuous learning capabilities, adopting a dynamic high-resolution strategy, and utilizing a high-quality bilingual dataset, InternVL 1.5 has demonstrated robust performance across a variety of multimodal benchmarks. Our evaluations indicate that the model achieves competitive performance with leading proprietary models, excelling particularly in OCR-related tasks and showing significant improvements in Chinese-related scene understanding. While InternVL 1.5 has contributed to the open-source multimodal understanding, the field continues to evolve with many challenges ahead. We aspire to further enhance InternVL’s capabilities and invite collaboration with the global research community, hoping to enrich and expand the reach of open-source models together.

### v2

@@ Abstract
In this report, we introduce InternVL 1.5, an open-source multimodal large language model (MLLM) to bridge the capability gap between open-source and proprietary commercial models in multimodal understanding. We introduce three simple improvements: (1) Strong Vision Encoder: we explored a continuous learning strategy for the large-scale vision foundation model—InternViT-6B, boosting its visual understanding capabilities, and making it can be transferred and reused in different LLMs. (2) Dynamic High-Resolution: we divide images into tiles ranging from 1 to 40 of 448$\times$448 pixels according to the aspect ratio and resolution of the input images, which supports up to 4K resolution input. (3) High-Quality Bilingual Dataset: we carefully collected a high-quality bilingual dataset that covers common scenes, document images, and annotated them with English and Chinese question-answer pairs, significantly enhancing performance in OCR- and Chinese-related tasks. We evaluate InternVL 1.5 through a series of benchmarks and comparative studies. Compared to both open-source and proprietary commercial models, InternVL 1.5 shows competitive performance, achieving state-of-the-art results in 8 of 18 multimodal benchmarks.
@@ 1 Introduction
We evaluated InternVL 1.5 on 18 representative multimodal benchmarks, which are categorized into four specific groups: OCR-related, general multimodal, mathematical, and multi-turn conversation benchmarks. Compared to both open-source and proprietary models, InternVL 1.5 shows competitive performance, achieving state-of-the-art results in 8 of 18 benchmarks. Notably, as shown in Figure 1, it even surpasses leading proprietary models like Grok-1.5V [125], GPT-4V [87], Claude-3 Opus [3], and Gemini Pro 1.5 [92] in four specific benchmarks, particularly in OCR-related datasets such as TextVQA [100], ChartQA [81], and DocVQA [82]. This evaluation indicates that InternVL 1.5 has effectively narrowed the gap between open-source models and leading commercial models. We hope that our approach and open-source model weights can contribute to the development of the MLLM community.
@@ 2.2 Open-Source MLLMs
The development of open-source MLLMs [138, 139, 124, 103, 2, 43, 48, 55, 51, 56, 69, 70, 110, 118, 120, 13] has significantly influenced the AGI landscape by integrating and enhancing capabilities in processing both visual and textual data. Over the past year, many open-source MLLMs have become well-known, including the LLaVA series [63, 62, 64], MiniGPT-4 [142], VisionLLM [116], Qwen-VL [5], CogVLM [117], Shikra [15], and others [119, 90, 18, 23]. However, these models are typically trained on images with small, fixed resolutions such as 336$\times$336, or 448$\times$448, which leads to sub-optimal performance on images with unusual aspect ratios or document data. To address this issue, many approaches have been explored for training on high-resolution images. Currently, there are two common technical routes: one involves designing a dual-branch image encoder [32, 77, 121, 76, 53], and the other involves dividing a high-resolution image into many low-resolution tiles [33, 47, 55, 57, 64, 68, 126, 127, 24]. Despite these explorations in high-resolution training, these open-source models still exhibit significant gaps in understanding documents, charts, and infographics, as well as recognizing scene texts, compared to leading commercial models.
@@ 3.4 High-Quality Bilingual Dataset
Pre-training Dataset. The pre-training dataset utilized in our InternVL 1.5 encompasses a diverse range of publicly accessible sources. We provide an overview of these datasets in Table 1(a). These datasets span multiple tasks, including captioning, which predominantly uses datasets such as Laion-EN [93], Laion-ZH [93], COYO [10], and GRIT [90], constituting 53.9% of the total data. Detection and grounding tasks utilize datasets like Objects365 [97], GRIT [90], and All-Seeing [119], making up 5.2%. For OCR tasks, we utilized large-scale datasets such as Wukong-OCR, LaionCOCO-OCR, and Common Crawl PDFs, which constitute 32.0% of our data. These datasets were constructed using PaddleOCR [49] to perform OCR on Chinese images from Wukong [29] and on English images from LaionCOCO [94]. Smaller OCR datasets include MMC-Inst [61], LSVT [105], ST-VQA [9], RCTW-17 [98], ArT [19], and others, accounting for 8.9% of the data, which focus on more specific or constrained OCR challenges. This diverse dataset assembly ensures robust model pre-training of InternVL, catering to varied linguistic and visual elements across tasks.
@@
XComposer2 [23] 8B 55.7 50.0
BLIP-2-XXL [50] 12B 54.8 49.1
Yi-VL-34B [130] 35B 54.2 48.6
Monkey-Chat [107] 10B 53.4 46.0
DeepSeek-VL [71] 7B 53.2 46.5
CogVLM-Chat [117] 17B 51.6 44.2
@@ 4.2.1 Quantitative Results on 18 Benchmarks
Figure 7: Examples for General QA. Blue highlights the standout answers, while red highlights incorrect responses.
@@ 4.3.1 Qualitative Results on Different Scenes
General QA. To compare the general capabilities of InternVL 1.5 and GPT-4V, we first conducted an experiment involving simple user queries with images requiring general knowledge. As shown on the left side of Figure 7, both two models respond accurately to the query, showcasing their proficiency in general topics. As shown on the right side of Figure 7, GPT-4V may excessively refuse to answer some questions due to its involvement in personal privacy.
@@
   Refer to caption User: Describe this image in detail. GPT-4V: The image shows a signboard with Chinese characters on it. From the text, we can deduce that the sign is likely for a business. The sign features a design that includes a martini glass with a flamingo, indicating that the establishment might serve drinks. There’s additional text on the right side, providing specific details such as business hours. It appears to be a nighttime business, likely a bar or restaurant, operating from 16:30 to 22:30. The other text provides further information about the services or specialties of the place. Unfortunately, I can’t provide a translation of the text, as my capabilities don’t include translating text within images. InternVL 1.5: The image shows a signboard with Chinese text, indicating a business establishment. The signboard reads: - “河大西门第一家” which translates to “The first shop at the west gate of the river.” - “正宗 老字号 五香味” which means “Genuine, old brand, five-spice flavor.” - “曹记焖烤鸡腿” which translates to “Cao’s roasted and baked chicken legs.” - “营业时间 16:30-22:30” indicating the business hours from 4:30 PM to 10:30 PM. - “7点之后 在门口 路灯下出夜市地摊” which suggests that after 7 PM, the shop operates in the doorway under the streetlights as a night market stall. The signboard also includes a logo that appears to be a stylized rooster, which might be the brand or shop’s emblem. The overall style of the signboard is quite traditional and straightforward, commonly seen in Chinese street markets or small businesses.
@@
Figure 8: Examples for OCR-related QA. Blue highlights the standout answers, while red highlights incorrect responses.
@@
Figure 11: Examples for Object Localization. Blue highlights the standout answers, while red highlights incorrect responses.
   Refer to caption User: What are the similarities and differences between these two pictures? GPT-4V: The similarities between the two images are that they both feature red pandas, a species known for their reddish-brown fur, bushy tails, and adorable faces. In both images, the red pandas are resting on tree branches with green foliage in the background, highlighting their arboreal (tree-dwelling) nature. As for the differences, the first image shows a red panda in a curled-up position, appearing to be asleep with its tail wrapped around its body for warmth. This posture also serves as camouflage, making the panda resemble a tree branch or part of the foliage. The angle of the shot is from above the animal, and the focus is on the texture and patterns of the fur. In contrast, the second image captures the red panda in a more alert and awake state, with its head raised and looking toward the camera. This posture allows a clear view of the panda’s facial features, such as its ears, eyes, and the white markings. The angle of the second image is more direct and at the same level as the animal, which gives a better sense of the panda’s overall appearance and stance. InternVL 1.5: The two pictures both feature red pandas, which are small, arboreal mammals native to the temperate forests of the Himalayas, southwestern China, and the northern part of Myanmar. They are known for their distinctive reddish-brown fur, white markings around their eyes, and bushy tails. In the first picture, the red panda is curled up and appears to be sleeping or resting on a branch. The background is blurred with green foliage, suggesting a natural, forested habitat. The focus is on the red panda’s relaxed posture and the texture of its fur. In the second picture, the red panda is sitting upright on a branch, with its front paws resting on the branch. The animal is looking directly at the camera, giving a more alert and curious expression. The background is also green and blurred, indicating a similar natural setting. The similarities between the two pictures include the presence of the red panda, the branch it is on, and the green, forested background. The differences lie in the red panda’s posture and expression, with the first picture showing a more relaxed and the second a more alert state. Additionally, the first picture has a more focused view of the red panda’s fur and form, while the second picture captures the animal’s face and eyes more clearly.
Figure 12: Examples for Multi-Image Dialogue. Blue highlights the standout answers.
@@
Object Localization. Evaluating machine learning models for their proficiency in object localization tasks is essential, especially in applications requiring precise spatial awareness. In our comparative analysis, the performance of the InternVL 1.5 model was juxtaposed with GPT-4V, focusing on their ability to accurately detect and localize objects within various environments. Our assessments ranged from simple object recognition in cluttered scenes to complex scenarios involving dynamic interactions among multiple entities. As illustrated in Figure  11, the results demonstrate that InternVL 1.5 not only localized objects with high accuracy but also exhibited a comparable understanding of spatial relationships, matching the performance of GPT-4V.
Multi-Image Dialogue. As shown in Figure 12, in this experiment, we ask InternVL 1.5 and GPT-4V to compare the similarities and differences between the two images. As can be seen, both GPT-4V and InternVL 1.5 provide detailed and accurate responses. Through this experiment, we discovered that although InternVL 1.5 was trained solely on single-image inputs, it exhibits strong zero-shot capabilities for multi-image dialogues.
@@ 5 Conclusion
This work introduced InternVL 1.5, an open-source MLLM designed to narrow the performance gap between open-source and proprietary models in multimodal understanding. By integrating a strong vision encoder with continuous learning capabilities, adopting a dynamic high-resolution strategy, and utilizing a high-quality bilingual dataset, InternVL 1.5 has demonstrated robust performance across a variety of benchmarks. Our evaluations indicate that the model achieves competitive performance with leading proprietary models, excelling particularly in OCR-related tasks and showing significant improvements in Chinese-related scene understanding. While InternVL 1.5 has contributed to the open-source multimodal understanding, the field continues to evolve with many challenges ahead. We aspire to further enhance InternVL’s capabilities and invite collaboration with the global research community, hoping to enrich and expand the reach of open-source models together.